OpenStudy (anonymous):

If the range of the function f(x) = 4 − 4x is {-20, -16, -8, 0, 4}, what is its domain? A) {0, 1, 2, 4, 6} B) {2, 3, 5, 6, 7} C) {1, 3, 5, 7, 9} D) {0, 1, 3, 5, 6}

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Im thinking you plug in each range or 'y' value and get its corresponding x value, then put them all together for the domain. so for the first one -20=4-4x x=?

OpenStudy (anonymous):

How do I figure out 'x'? Sorry, really bad at this, which is why I'm asking for assistance xD.

OpenStudy (anonymous):

Alrighty, just solve for x like you would any equation subtract 4 from both sides -20-4=-4x -24=4x divide both sides by -4 -24/-4=x 6=x so thats one value for the domain, next one you need to solve -16=4-4x

OpenStudy (anonymous):

-20=4-4x -16=4-4x -8=4-4x 0=4-4x 4=4-4x solve each and find the value of x.

OpenStudy (anonymous):

Okay, so now I would go like so.. -16=4-4x Now subtract 4 from both sides, which would bring in result of -16-4=4x -20=4x Then divide both sides by 4 and that would be -20/4=4x/4 -5=x Correct?

OpenStudy (anonymous):

Close, forgot the negative sign on -4x -20/-4=-4x/-4 5=x

OpenStudy (anonymous):

now for -8=4-4x

OpenStudy (anonymous):

Ok. -8=4-4x -8-4=-4x -12=-4x -12/-4=-4x/-4 Which in final results is 3=x. Right?

OpenStudy (anonymous):

You've got it. next one 0=4-4x and then 4=4-4x

OpenStudy (anonymous):

0=4-4x 0-4=-4x -4=-4x 1=x, right? for the first one.. Then 4=4-4x 4-4=-4x 0=-4x 0=x, correct?

OpenStudy (anonymous):

Yep, now put all of our answers together from smallest to largest.. {0,1,3,5,6}

OpenStudy (anonymous):

Thank you so much, you were an extremely big help! :) This was confusing the heck outta me. Now thanks to you, I got it! THANKS SO MUCH! (:

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

correction -16=4-4x -16-4=-4x -20=-4x \[\frac{ -20 }{-4 }=x or x=5\] similarly for others